# Learning from Imperfect Human Feedback: a Tale from Corruption-Robust $\text{Dueling}^{*}$

### v1

@@ 4 The Intrinsic Limit of Learning from Imperfect Human Feedback
Assume that the action space $\mathcal{A}$ is contained in a $d$-dimensional unit ball. Consider the utility function $\mu$ in the form $\mu_{\theta}(a)=\theta^{\top}a-\frac{1}{2}\|a\|^{2}_{2}$, where $\theta\in\mathbb{R}^{d},\|\theta\|_{2}\leq 1$, is a random vector. Under corruption induced by $\rho$-imperfect user (Def. 1), for any fixed number of rounds $T$ and $d\leq\frac{1}{C_{\kappa}}T^{1-\rho}$, there exists a preference parameter $\theta$ such that for any learner $\mathcal{L}$ under reward feedback, even with the knowledge of $\rho$, has to suffer regret $\text{Reg}_{T}:=\mathbb{E}\{\sum^{T}_{t=1}\mu_{\theta}(a^{*})-\mu_{\theta}(a_{t})\}\geq\frac{d}{4}C_{\kappa}T^{\rho}.$
Proof (Proof Sketch of Lemma 1). We aim to show that there exists a corruption strategy such that the regret incurred by any learner $\mathcal{L}$ with corrupted reward feedback is no less than $\Omega(dT^{\rho})$. The formulation of such a corruption strategy hinges on the observation that the regret incurred by $\mathcal{L}$ is bounded below by the sum of the Kullback-Leibler (KL) divergence between the distribution of the corrupted reward feedback conditioned on different possible values of $\theta$. Specifically, let $\theta$ be uniformly drawn from $\{-\beta,\beta\}^{d}$. Given an action $a_{t}$, conditioned on $\theta_{i}>0$, the corrupted reward feedback $\hat{v}_{t}$ is
$\hat{v}_{t}=\mu_{\theta}(a_{t})+c_{t}(a_{t}|\theta_{i}>0)+\xi_{a_{t}}=\underbrace{\left(-\frac{1}{2}\|a_{t}\|^{2}+\sum_{j\neq i}\theta_{j}a_{t,j}\right)+\beta a_{t,i}+c_{t}(a_{t}|\theta_{i}>0)}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mu_{1}}}+\xi_{a_{t}}.$ (2)
Conditioned on $\theta_{i}<0$, the corrupted reward feedback is
$\hat{v}^{\prime}_{t}=\mu_{\theta}(a_{t})+c_{t}(a_{t}|\theta_{i}<0)+\xi_{a_{t}}=\underbrace{\left(-\frac{1}{2}\|a_{t}\|^{2}+\sum_{j\neq i}\theta_{j}a_{t,j}\right)-\beta a_{t,i}+c_{t}(a_{t}|\theta_{i}<0)}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mu_{2}}}+\xi_{a_{t}}.$ (3)
Consider the following corruption strategy which sets $c_{t}(a_{t}|\theta_{i}>0)=-\beta a_{t,i}$ and $c_{t}(a_{t})=\beta a_{t,i}$ to minimizes the KL divergence between $\hat{v}_{t}$ and $\hat{v}^{\prime}_{t}$. Together with $1$-strongly concavity of $\mu_{\theta}$, we can establish
$\mathbb{E}\left\{\sum^{T}_{t=1}\mu_{\theta}(a^{*})-\mu_{\theta}(a_{t})\right\}\geq\frac{1}{2}\max\left\{\sum^{T}_{t=1}(\frac{C_{\kappa}T^{\rho-1}}{\beta}-\beta)^{2},\frac{dT\beta^{2}}{2}\right\}.$ (4)
Since $d\leq\frac{1}{C_{\kappa}}T^{1-\rho}$, we can set $\beta=\sqrt{C_{\kappa}}T^{\frac{\rho-1}{2}}$ to satisfy $\|\theta\|_{2}\leq 1$ and to optimize the lower bound in Equation (4). Then we obtain $\mathbb{E}\left\{\sum^{T}_{t=1}\mu_{\theta}(a^{*})-\mu_{\theta}(a_{t})\right\}\geq dC_{\kappa}T^{\rho}/4$. Notice that the corruption level each round is bounded by $\beta\|a_{t}\|_{\infty}$. In the scenario when the radius of the action space $\mathcal{A}$ is upper bounded by $\beta$, the corruption budget each round is bounded by $C_{\kappa}t^{\rho-1}$. It implies that the corruption is induced by $\rho$-imperfect user, completing the proof.
5 The Efficiency-Robustness Tradeoff in LIHF
Section 4 shows that a lower bound regret of $\Omega(d\max\{\sqrt{T},T^{\rho}\})$ is intrinsic for learning from $\rho$-imperfect user, even with known $\rho$. In this section, we study what would be possible — but in a significantly more challenging learning setting with arbitrary and agnostic corruption. The reason for fighting for such generality is due to practical considerations; in reality, it is rarely possible to truly predict users’ irrationality or behaviors. Technique-wise, gradient-based methods are the most commonly used for addressing continuous action spaces, and so are our approaches. The key novelty in our results is not about the development of fundamentally new algorithmic techniques, but rather the conduct of more fine-grained analysis about two popular gradient-based methods — Dueling Bandit Gradient Descent (DBGD (Yue and Joachims, 2009)) and Noisy Comparison-based Stochastic Mirror Descent (NC-SMD) (Kumagai, 2017) — and the proof of how these algorithms can be “tailored” to achieve robustness against corruption. On a conceptual level, our key insight is that one can carefully tune the learning rates of these gradient-based methods to gracefully trade off learning efficiency for adversarial robustness — a interesting concept which we coin the efficiency-robustness tradeoff. Notably, en route developing our result, we also pin down a tight lower bound of $\Omega(T^{3/4})$ for the well-known DBGD algorithm which surprisingly remained an open question to date despite extensive studies of this algorithm.
Our finding reveals the intrinsic tradeoff between robustness and efficiency for gradient-based algorithms under agnostic corruption (Theorem 2). In general, the tradeoff is controlled by the exploration size in Algorithm 1 (depends on $\alpha$) (experimentally verified at Appendix D.1.4). A higher tolerance of agnostic corruption comes at a cost of worse regret in scenario without corruption, shown by Figure 1.
1 for $t\in[T]$ do
2       Sample unit vector $u_{t}$ uniformly and set $a^{\prime}_{t}:=\mathcal{P}_{\mathcal{A}}\left(a_{t}+\delta u_{t}\right)$.
4       Compute gradient $\hat{g}_{t}:=-\frac{d}{\delta}\hat{\mathcal{F}}(a^{\prime}_{t},a_{t})u_{t}$.
5       Set step size $\eta:=\frac{\gamma\delta}{d}$ and update $a_{t+1}:=\mathcal{P}_{\mathcal{A}}(a_{t}-\eta\hat{g}_{t})$.
Algorithm 1 Dueling Bandit Gradient Descent (DBGD)
Theorem 2 (Efficiency-Robustness Tradeoff in DBGD).
For any $\alpha\in(0,\frac{1}{4}]$ and any number of round $T$ satisfying $T>\left(\frac{\sqrt{2Rd}L_{\mu}L_{2}}{\sqrt{13L}L_{\sigma}}\right)^{4}$, Algorithm 1 parameterized with $\gamma=\frac{R}{\sqrt{T}}$ and $\delta=\frac{\sqrt{2Rd}}{\sqrt{13L}T^{\alpha}}$ achieves $\text{Reg}_{T}\leq O(\sqrt{d}T^{1-\alpha}+\sqrt{d}T^{\alpha}C)$ for any total corruption level $C$.
At the core of this proof is a novel development of a regret decomposition lemma (Lemma 2), where we decompose the regret into two parts: regret of decision making under uncertainty and regret due to feedback imperfection. This decomposition has been proven to be useful for controlling regret of interactive decision-making in reinforcement learning Foster et al. (2023). However, to the best of our knowledge, Lemma 2 is the first to exhibit such a decomposition for dueling feedback, which contains much sparser information than the direct reward feedback as in RL or classic online learning in the aforementioned studies. Hence we believe the lemma may be of independent interest for future research on dueling bandits under corruption or erroneous observation. Our proof employs new techniques and is also more involved. Previous proofs of the decomposition lemma for reward feedback crucially hinges on an online estimation oracle Foster et al. (2023), which however cannot be constructed in our problem with dueling feedback with agnostic corruption. We thus have to develop new proof ideas tailored specifically for dueling bandit feedback. We start from quantifying the bias $b_{t}$ caused by adversarial corruption in gradient estimation, and then show that we can control the cumulative impact of bias by the exploration size $\delta$, i.e. $\sum^{T}_{t=1}\mathbb{E}\left(b_{t}^{\top}(a_{t}-a^{*})\right)\leq 2RdL_{\sigma}C/\delta$. Consequently, Lemma 2 shows that the regret of decision and observation error are jointly controlled by the exploration size $\delta$. We choose $\delta$ in order of $\Theta(T^{-\alpha}),\alpha\in(0,1/4]$ to achieve a balance between optimizing the learning objective (1) and controlling the gradient bias arising from the corrupted dueling feedback. Following this, we show $\text{Reg}_{T}$ is upper bounded by $O(\sqrt{d}T^{1-\alpha}+\sqrt{d}T^{\alpha}C)$.
Figure 1: Efficiency-Robustness Tradeoff
Lemma 2 (Regret Decomposition).
Let $b_{t}:=\frac{d}{\delta}\left(\hat{\mathbb{P}}(a^{\prime}_{t}\succ a_{t})-\mathbb{P}(a^{\prime}_{t}\succ a_{t})\right)u_{t},u_{t}\sim\mathbb{S}$, and choose $\lambda=\frac{L_{\sigma}}{L_{\sigma}-\delta L_{\mu}L_{\sigma}}$, we have
$\text{Reg}_{T}$ $\leq\underbrace{\lambda\left(\frac{R^{2}d}{\gamma\delta}+\frac{T\gamma d}{\delta}+13\delta LT\right)}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Regret\hskip 2.84526ptof\hskip 2.84526ptDecision}}+\underbrace{2\lambda\sum^{T}_{t=1}\mathbb{E}\left(b_{t}^{\top}(a_{t}-a^{*})\right)}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}Observation\hskip 2.84526ptError}}.$
Note that Algorithm 1 works for general concave user utility. However, if the user displays a strongly concave utility, we can tune learning rate of NC-SMD to get better tradeoff. Specifically, we can get better regret upper bound $\tilde{O}(d\sqrt{T})$ in scenario without corruption at the same time affording $O(T^{3/4})$ agnostic corruption (see Prop 2). The proof of Proposition 2 shares similar structure as the proof Theorem 2, however working with a different algorithm. We direct readers to Appendix C.1 and C.4 to see full statement of Theorem 2, details of the Algorithm NC-SMD and proof of Proposition 2.
Proposition 2 (Efficiency-Robustness Tradeoff in NC-SMD).
@@ 5 The Efficiency-Robustness Tradeoff in LIHF
For any $\alpha\in[\frac{1}{2},1)$, if the utility function is strongly concave, by choosing learning rate $\eta=\frac{\sqrt{\log T}}{2d}T^{-\alpha}$, for the Algorithm NC-SMD, we have $\text{Reg}_{T}\leq\tilde{O}(dT^{\alpha}+\sqrt{d}T^{\frac{1}{2}(1-\alpha)}C+dC)$ for any total corruption level $C$.
Figure 2: DBGD Lower Bound
An interesting corollary of Theorem 2 is that we can use it to show the DBGD upper bound proved in (Yue and Joachims, 2009) is tight, which to the best of our knowledge, has not been realized in the previous literature. In the following we provide a sketch of Corollary 1. It has been verified by experiments (see Figure 2, simulation details in Appendix D.1.1). We direct readers to Appendix C.2 for proof details.
Corollary 1 (Minimax Lower Bound for DBGD).
@@
There exists a linear utility function $\mu$ and a link function $\sigma$ such that choosing $\alpha=1/4$, $\text{Reg}_{T}$ suffered by Algorithm 1 is at least $\Omega\left(T^{3/4}\right)$.
Proof  (Proof Sketch of Corollary 1). Corollary 1 shows that the standard DBGD algorithm with $\alpha=\frac{1}{4}$ can afford $O(T^{3/4})$ agnostic corruption. Moreover, we observe that we can make the sub-optimal arm and the optimal arm indistinguishable by using the total corruption level in the same order of the number of times sub-optimal arm that has been pulled. Therefore, there must exist an problem instance such that the sub-optimal arm has been pulled at least $\Theta(T^{3/4})$ times by DBGD. If not, we can create the parallel world to make DBGD suffer linear regret in one of the environments by using total corruption level less than $\Theta(T^{3/4})$, which contradicts Theorem 2. Intuitively, it says that if an environment is ”easy” to attack, the agnostic corruption level which an algorithm could afford can not exceed the worst possible regret suffered by the algorithm in scenario without corruption.
@@ 6 Experiments
In this section, we study the empirical performance of DBGD and NC-SMD to validate our theoretical analysis by running simulations on both synthetic datasets and real-world recommendation data in comparison with several baseline algorithms. Experimental details can be found at Appendix D.
6.1 Synthetic Data
@@
Experiment Setup. We consider a standard experiment setup, which adopts a strongly concave utility $\mu_{\theta}(a):=\theta^{\top}a-\frac{1}{2}\|a\|^{2}_{2}$, a logistic link function $\sigma(x)=\frac{1}{1+\exp(-x)}$. We choose $d=5$, and $T=10^{5}$. Our action space $\mathcal{A}$ is a $d$-dimensional ball with radius $R=10$. The preference parameter $\theta$ is randomly sampled from the surface of $\mathcal{A}$. In our problem setting, the optimal action $a^{*}$ is $\theta$ and $\mu_{\theta}(a^{*})=50$. We simulate agnostic arbitrary corruption for $C=T^{\rho},\rho\in[0.5,0.75]$. Specifically, we force the user to submit her least preferred item to the algorithm over the first $C$ rounds (for experiments on corruption induced by $\rho$-imperfect user, see Appendix D.1.3). For each value of $\rho$, we repeat the experiments for $5$ times, each under a different seed. We use Doubler and Sparring proposed by Ailon et al. (2014) as the baseline algorithms for comparison. We set bandit gradient descent (Flaxman et al., 2004) as the black-box bandit algorithm.
Figure 3: We show the log-log plot for regret occurred by different algorithms under arbitrary agnostic corruption. For each $\rho$, we present the line plot of average regret over $5$ simulations. The shaded region is the average regret $\pm$ one standard deviation. In addition, we fit the order of the average regret, i.e. $\mathbb{E}(\text{Reg}_{T})=T^{o}$, by least squares using last $1\%$ of data.
Results and Discussion. Figure 3 shows that standard DBGD ($\alpha=1/4$) and NC-SMD ($\alpha=1/2$) are able to tolerate $O(T^{3/4})$ unknown corruption, aligning with the theoretical result. Experimentally, NC-SMD enjoys the best performance under strongly concave utility. Empirically, Sparring has comparable performance as DBGD, even though it is a heuristic approach with no theoretical guarantee. The worst performance of Doubler might be attributed to the fact that its theoretical guarantee is only applicable to linear link function, which is different from our experiment setup.
6.2 Spotify Recommendation Data
Figure 4: Results on Spotify Recommendation
Evaluation Setup. We also evaluate our approach on Spotify recommendation data (Spotify, 2020). The objective is to recommend songs to incoming users. This dataset includes $17\times 10^{4}$ songs ($|\mathcal{A}|=17\times 10^{4}$), each is described by 15 distinct features (d = 15). We create 5 different user types by clustering the dataset into 5 groups. We use the average of the song embedding within each group as the user’s preference vector. Our aim is to recommend songs to users that have the highest cosine similarity (utility function) with their preference vectors. We rescale the cosine similarity to $[-100,100]$. We simulate corruption induced by $\rho$-imperfect user to model imperfect human feedback. We only test the performance of DBGD and Sparring because NC-SMD only works under strongly concave utility and Doubler only works under linear link function.
Results and Discussion. Figure 4 show that the performance of DBGD aligns with theoretical prediction on real-world recommendation data, even in the scenario when $\mathcal{A}$ is discrete and nonconvex. Both DBGD and Sparring are robust to agnostic corruption induced by $\rho$-imperfect user, however, empirically it seems that Sparring is invariant to $\rho$. This might because the song embeddings are highly sparse, which results in the fact that the proposed arms according to the heuristic approach are easy to distinguish, even under agnostic corruption.
Additional Experiments. We also test the performance of Versatile-DB (Saha and Gaillard, 2022), an approach robust to agnostic corruption with theoretical guarantee. However, this approach is not computationally efficient. When $K=2000$, running 100 iterations takes around 20 CPU hours and it’s performance is dominated by DBGD. We direct readers to Appendix D.2.1 to see the details.

### v2

@@ 7       Set $a_{t+1}=\nabla\mathcal{R}_{t}^{-1}(\nabla\mathcal{R}_{t}(a_{t})-\eta_{\rho}\hat{g}_{t})$
If $A=\mathbb{E}\left(\sum^{T}_{t=1}[\mu(a^{*})-\mu(a_{t})]\right)\leq\mathcal{O}(T^{\frac{1}{2}+\psi})$ for some $\psi\in[0,\frac{1}{2})$, then we must have $B=\mathbb{E}\left(\sum^{T}_{t=1}t^{\rho-\frac{1}{2}}\sqrt{\mu(a^{*})-\mu(a_{t})}\right)\leq\mathcal{O}(T^{-\frac{1}{4}+\frac{\psi}{2}+\rho})$.
To prove Lemma 5, we first bound term $B$ by $\sum^{T}_{t=1}t^{\rho-1/2}\sqrt{\mathbb{E}[\mu(a^{*})-\mu(a_{t})]}$ using the concavity of the square root function. The key novelty of our proof is to use the Abel’s equation (Williams, 1963) to re-arrange the term $\sum^{T}_{t=1}t^{\rho-1/2}\sqrt{\mathbb{E}[\mu(a^{*})-\mu(a_{t})]}$ and establish its connection with another term $C=\sum^{T}_{t=1}\sqrt{\mathbb{E}[\mu(a^{*})-\mu(a_{t})]}$. We finally connect term $C$ with the term $A$ in the lemma statement, by proving $\sum^{T}_{t=1}\sqrt{\mathbb{E}[\mu(a^{*})-\mu(a_{t})]}\leq\mathcal{O}(T^{\frac{3}{4}+\frac{\psi}{2}})$ if $\mathbb{E}\left(\sum^{T}_{t=1}\mu(a^{*})-\mu(a_{t})\right)\leq\mathcal{O}(T^{\frac{1}{2}+\psi})$ via concavity of square root. Together with the decay scale $t^{\rho-1}$, which allows $t^{\rho-1}-(t+1)^{\rho-1}\leq t^{\rho-2}$, we obtain the tight bound $\mathcal{O}(T^{-\frac{1}{4}+\frac{\psi}{2}+\rho})$. [^5] If corruption $c_{t}$ was arbitrary, then $t^{\rho-1/2}$ is $\sqrt{t}|c_{t}|$, and the bound could deteriorate to $\mathcal{O}(T^{\frac{2\rho}{3}})$ in the first $T^{\rho}$ rounds.
Armed with Lemma 5, we can use it to prove the induction claim (Lemma 6). The challenge in this proof lies in identifying the constant $K$, which must hold across all iterations, $k$, of the analysis to ensure the bound remains stable and does not explode.
Lemma 6 (Induction Claim).
Consider $T>\sqrt{2L^{\mu}L^{\sigma}_{1}R}$ and $\eta_{\rho}=\frac{\sqrt{\log T}}{dT^{\max\{1/2,\rho\}}}$. If $\text{Reg}_{T}\leq 144dK\sqrt{\log T}T^{\rho+\frac{\frac{3}{2}-\rho}{2^{k}}}$ for some integer $k$, then we must also have $\text{Reg}_{T}\leq 144dK\sqrt{\log T}T^{\rho+\frac{\frac{3}{2}-\rho}{2^{k+1}}}$. $K$ is a instance-dependent constant, where $K:=\max\left\{C_{0},\frac{C_{2}C_{\kappa}}{\sqrt{\log T}},(L^{\sigma}C_{\kappa})^{2},2\sqrt{\lambda}RC_{\kappa}L^{\sigma}\right\}$, $C_{0}$ and $C_{2}$ defined in Lemma 3 and 4 respectively,
4.1 Additional Applications of the Above Regret Analysis Framework
To showcase useful applications of the new regret analysis framework above for proving Theorem 2, in this subsection we employ the framework to study continuous dueling bandits under arbitrary and agnostic corruption — i.e., remove decaying constraints on $c_{t}$ and $\rho$ is unknown to the learner. We analyze natural variants of RoSIMD, which apply to strongly concave utilities, and the well-known algorithm Dueling Bandit Gradient Descent (DBGD) (Yue and Joachims, 2009), which apply to general concave utilities. The proofs of these results are direct applications of the analysis from Steps 1-3 outlined above (without Step 4, since it is the only one requiring knowledge of $\rho$ and decaying corruption). Additionally, we present these results to offer a set of principled benchmark algorithms for the experiments in Section 5.
Proposition 2 (Efficiency-Robustness Tradeoff in RoSMID).
@@ 4.1 Additional Applications of the Above Regret Analysis Framework
If the total corruption level satisfies $\sum^{T}_{t=1}|c_{t}(a_{t},a^{\prime}_{t})|\leq\mathcal{O}(T^{\rho})$, then for any $\alpha\in[\frac{1}{2},1)$, if the utility function is strongly concave, for a sufficiently large round $T$, by choosing learning rate $\eta_{\rho}=\frac{\sqrt{\log T}}{2d}T^{-\alpha}$
  1. 1.
    (Robustness) RoSMID incurs $\text{Reg}_{T}\leq\tilde{\mathcal{O}}(dT^{\alpha}+\sqrt{d}T^{\frac{1}{2}(1-\alpha)+\rho}+dT^{\rho})$.
  2. 2.
@@
    (Efficiency) There exists a strongly concave utility function $\mu$ and a link function $\sigma$ such that $\text{Reg}_{T}$ suffered by RoSMID is at least $\Omega\left(T^{\alpha}\right)$ in scenario without corruption.
Proposition 3 (Efficiency-Robustness Tradeoff in DBGD).
If the total corruption level satisfies $\sum^{T}_{t=1}|c_{t}(a_{t},a^{\prime}_{t})|\leq\mathcal{O}(T^{\rho})$, then for any $\alpha\in(0,\frac{1}{4}]$, if utility function $\mu$ is generally concave, for a sufficiently large round $T$, choosing $\gamma=\frac{R}{\sqrt{T}}$ and $\delta=\frac{\sqrt{2Rd}}{\sqrt{13L^{\sigma}L^{\mu}}T^{\alpha}}$ for DBGD
  1. 1.
    (Robustness) $\text{Reg}_{T}$ incurred by DBGD satisfies $\text{Reg}_{T}\leq\mathcal{O}(\sqrt{d}T^{1-\alpha}+\sqrt{d}T^{\alpha+\rho})$.
  2. 2.
    (Efficiency) There exists a linear utility function $\mu$ and a link function $\sigma$ such that $\text{Reg}_{T}$ suffered by DBGD is at least $\Omega\left(T^{1-\alpha}\right)$ in scenario without corruption.
Both propositions illustrate the tradeoff between learning efficiency and robustness to adversarial corruption, achieved through their tunable learning rate $\eta_{\rho},\gamma,\delta$ (learning rate in DBGD is $\frac{\gamma\delta}{d}$; we direct reader to Algorithm 3 for detail). Specifically, greater tolerance for agnostic corruption (i.e., larger $\alpha$ in Proposition 2) results in worse regret in non-corrupt settings. A similar robustness-accuracy trade-off has been studied in classification problems with adversarial examples, both empirically (Tsipras et al., 2018) and theoretically (Zhang et al., 2019). However, to our knowledge, this is the first formal analysis of such a trade-off in online learning. While expanding the confidence region for UCB-type algorithms to handle agnostic corruption in linear contextual dueling bandits increases tolerance for corruption but worsens regret—sharing a similar idea of a trade-off (Di et al., 2023)—it’s unclear if this approach necessarily reduces efficiency. In contrast, our work demonstrates that adjusting learning rates in gradient-based algorithms to enhance robustness against agnostic corruption inevitably decreases efficiency when no corruption is present. In particular, the efficiency statements in both propositions can be interpreted as algorithm-dependent lower bounds, concretely demonstrating that for certain problem instances, increasing robustness necessarily decreases the learning efficiency of the algorithm.
@@ 5 Experiments
In this section, we validate our theoretical analysis of RoSMID and DBGD under $\rho$-Imperfect Human Feedback through simulations. We also compare their performance against two baseline algorithms: Doubler and the heuristic algorithm, Sparring, proposed by Ailon et al. (2014). For Sparring, we use bandit gradient descent (Flaxman et al., 2005) as the black-box bandit algorithm. We direct readers to Appendix E for implementation details.
@@
Experiment Setup. We consider a standard experiment setup, which adopts a strongly concave utility $\mu_{\theta}(a):=\theta^{\top}a-\frac{1}{2}\|a\|^{2}_{2}$, and a logistic link function $\sigma(x)=\frac{1}{1+\exp(-x)}$. We choose $d=5$, and $T=10^{5}$. Our action space $\mathcal{A}$ is a $d$-dimensional ball with radius $R=10$. The preference parameter $\theta$ is randomly sampled from the surface of $\mathcal{A}$. In our problem setting, the optimal action $a^{*}=\theta$ and $\mu_{\theta}(a^{*})=50$. We simulate $\rho$-Imperfect Human Feedback for $\rho\in[0.5,1]$.
Figure 1: Robustness of RoSMID for $\rho$-LIHF
Results and Discussion. Figure 1 presents a log-log plot of regret versus iteration, $t$, which the order of regret in $T$ is represented by the slope of the line. Each line represents the average over five trials with different random seeds, and each line corresponds to a distinct value of $\rho$. The shaded region indicates $\pm$ one standard deviation. In the legend, “o” denotes the estimated order of regret. We estimate it using least squares on the last 1% of the data. Figure 1 supports Theorem 2, as for $\rho\in[0.5,1)$, the estimated slope is less than $\rho$, suggesting that the choice of $\eta_{\rho}=\frac{\sqrt{\log T}}{dT^{\max\{1/2,\rho\}}}$ for RoSMID results in a regret bound of at most $\tilde{\mathcal{O}}(T^{\rho})$.
Figure 2 compares the performance of our proposed algorithms, variants of DBGD and RoSMID, with two baseline algorithms, Sparring and Doubler, in the setting when $c_{t}$ is arbitrary and agnostic. We set the parameter $\alpha=1/4$ for DBGD and $\alpha=1/2$ for RoSMID. When $\rho$ is unknown to the algorithms, these parameter choices enable DBGD and RoSMID to tolerate $\mathcal{O}(T^{3/4})$ levels of unknown corruption, since the slope estimates are less than 1, implying sublinear regret, aligning with our theoretical predictions. Experimentally, RoSMID demonstrates the best performance under strongly concave utility functions. Sparring performs comparably to RoSMID, despite being a heuristic approach without theoretical guarantees. In contrast, Doubler exhibits the worst performance, likely because its theoretical guarantees are only applicable to linear link functions, which differs from our experimental setup. We direct readers to Appendix E for experiment results with other choice of $\alpha$ for efficiency-robustness tradeoff.
Figure 2: For each algorithm, we tested its performance under $\rho$-Imperfect Human feedback with $\rho=0.5,0.6,0.75$. For each $\rho$, we presented a line plot of the average regret over five simulations, accompanied by $\pm$ one standard deviation shown by the shaded region. In the legend, $o$ denotes the estimated line slope, calculated using least squares on the last $1\%$ of the data.